OpenStudy (anonymous):

will give medal if help Find the experimental probability. Sheila drew single marbles from a bag, recorded the outcome, and replaced the marble. In 50 draws, Sheila drew 20 white marbles and 30 black marbles. What is the experimental probability that the next marble will be black? a 2/5 b 3/5 c 2/3 d 3/7

OpenStudy (lucaz):

30 black in 50 draws = 30/50 = 3/5 P(black)=3/5

OpenStudy (anonymous):

thanks